If a car accelerates from rest to 100 ms·1 in 5 seconds, what was itsacceleration?
1
Expert's answer
2020-04-02T11:01:13-0400

As per the given question,

Initial speed of the car (u)=0

Final speed of the car (v)=100 m/sec

time (t)= 5 sec

we know that

"v=u+at"

Now, substituting the value in equation,

"\Rightarrow 100=0+a\times 5"

"a=\dfrac{100}{5}=20 m/sec^2"